OpenStudy (anonymous):

Michael makes 90% of his free throws in basketball. If he shoots 20 free throws, and if his chance of making each one is independent of the other shots, what is the probability that he makes… a. all 20? b. exactly 18? c. at least 18?

hero (hero):

They must be talking about Michael Jordan

OpenStudy (anonymous):

Haha... I think that you missed the point of the question.

OpenStudy (anonymous):

Probability make the 20 free throws is \[ .9^{20} =0.121577 \]

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

Exactly 18 \[ \frac{ 20!}{18! 2!}\,0.9^{18}\, 0.1^2=0.28518 \]

OpenStudy (anonymous):

At Least 18 \[ \frac{ 20!}{19! 1!} 0.1^1 0.9^{19}+ \frac{ 20!}{18! 2!}0.1^2 0.9^{18}+0.9^{20}=0.676927 \]

OpenStudy (anonymous):

Did you understand what I did?

OpenStudy (anonymous):

Yeah. Thanks you soo much. I really appreciate it.

OpenStudy (anonymous):

yw